OpenStudy (anonymous):

What is the equation of the axis of symmetry for the function shown below? y-1= -2(x+3)^2

OpenStudy (anonymous):

axis of symmetry is x+3=0

OpenStudy (anonymous):

so 0?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

no ,x+3=0,or x=-3

OpenStudy (anonymous):

The general axis of symmetry formula for a parabola is x = -b/2a You should first separate y from everything else: y = -2(x+3)^2 + 1 and reach a point where you have y = ax^2 + bx + c So for your equation: y = -2(x^2 + 6x + 9) + 1 y = -2x^2 - 12x - 18 + 1 y = -2x^2 - 12x - 17 Then you would have a = -2, b = -12, c = -17 So the axis of symmetry: x = -(-12)/2*(-2) x = 12/-4 x = -3

OpenStudy (anonymous):

Oh, or just what surjithayer said :)

OpenStudy (anonymous):

\[-\frac{ 1 }{2 }\left( y-1 \right)=\left( x+3 \right)^2\] it is a downward parabola,with vertex at (-3,1) and equation of axis of symmetry is x+3=0 or x=-3 cogs is also correct but approach is different than me.